Segment AD has a midpoint B, BD has a midpoint C. If A has coordinates (-5, 4) and C has coordinates (10, 11), what are the coordinates of B and D?
Answer by stanbon(75887)

Segment AD has a midpoint B, BD has a midpoint C. If A has coordinates (-5, 4) and C has coordinates (10, 11), what are the coordinates of B and D?
---------------------------
That's a pretty messy problem, but doable.
----------------------------------------------
Draw the picture so you can follow what is happening.
----------------
Label point D as (a,b)
Label point B as (x,y)
-------------------------
EQUATIONS:
Since C is the midpoint of BD
x+a = 2*10
y+b = 2*11
-----------------
Since B is the midpoint of AD
a-5=2x
b+4=2y
-------------------
Look at these two equations:
x+a = 20
2x-a = -5
Add them to solve for x:
3x = 15
x = 5
Substitute to solve for "a":
5+a=20
a= 15
---------------------
Look at these two equations:
y+b = 22
2y-b=4
Add to solve for y:
3y = 26
y = 26/3
Substitute to solve for "b":
(26/3)+b = (66/3)
b = 40/3
------------------------
Coordinates:
(a,b) = (15,40/3)
(x,y) = (5,26/3)
===============================
Cheers,
Stan H.